# Cloud gap-filling with deep learning for improved grassland monitoring

### v1

@@ 7 Conclusion
The Sentinel-2 satellite missions provide frequent optical images for the continuous monitoring of grassland activity. However, relying on optical images alone for monitoring is restricted by cloud coverage. To overcome this limitation, we proposed a CNN-RNN deep learning architecture that makes use of all available cloud-free optical images (Sentinel-2), along with weather-independent SAR data (Sentinel-1), to create continuous NDVI time series for grassland fields, with a temporal resolution of 6 days. The model has been trained on six different regions of Lithuania to account for different cloud coverage profiles and agro-climatic conditions.
The results showed that our method performed better than alternative gap-filling techniques (linear, Akima, quadratic interpolation) when evaluated across various scenarios to demonstrate its transferability and generalization capacity. Additionally, we assessed the importance of the generated continuous NDVI time series for the downstream task of grassland mowing event detection. The outcomes suggest that our methodology significantly improves the detection accuracy by i) identifying events that would have otherwise gone undetected and ii) eliminating instances of false event predictions caused by cloud presence.
Finally, we investigated our model’s capacity to detect grassland mowing events. Overall, the outcomes are promising, with f1-score of 85%, surpassing the performance of the two baseline methodologies by more than 6%.
@@ Acknowledgements
Data availability
Data will be made available on request.

### v2

@@ 7 Conclusion
The Sentinel-2 satellite missions provide frequent optical images for the continuous monitoring of grassland activity. However, relying on optical images alone is often restricted by cloud coverage. To overcome this limitation, we proposed a CNN-RNN deep learning architecture that makes use of all available cloud-free optical images (Sentinel-2) and weather-independent SAR data (Sentinel-1), to produce continuous NDVI time series for grassland fields, with a temporal resolution of 6 days. The integration of cloud-free Sentinel-2 data into our model is a significant advancement of this work, enhancing monitoring capabilities compared to relying solely on SAR data. The model has been trained on six different regions of Lithuania to account for different cloud coverage profiles and agro-climatic conditions.
Our method outperformed alternative gap-filling techniques (linear, Akima, and quadratic interpolation) in various evaluation scenarios, demonstrating its transferability and generalization capacity. Most importantly, the generated continuous NDVI time series significantly enhanced the accuracy of mowing event detection by i) identifying events that would have otherwise gone undetected and ii) eliminating instances of false event predictions caused by clouds. Our model demonstrated a promising F1-score of 84%, surpassing the respective rates of baseline methodologies by more than 6%.
@@
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT in order to improve readability. After using this tool, the authors reviewed and edited the content as needed and took full responsibility for the content of the publication.
@@ Acknowledgements
The work of V. Sitokonstantinou has been partly supported by GVA PROMETEO AI4CS project on ‘AI for complex systems’ (2022-2026) - CIPROM/2021/056.
Availability of data and Materials